# XpertAI: uncovering model strategies for sub-manifolds

### v1

@@ Appendix B How to train and select good range experts?
Figure 10: Overview model performance for the regression benchmarks presented in Sec. 5.Left: original label distribution colour-coded by range expert. Centre: original label vs. original model output. Right: original model output vs. output of the respective range experts. Note, that the results shown are from range experts retrained to obtain a more disentangled representation in connection with LRP.
Figure 11: Overview of model performance on the rMNIST problem for 3, 6 and 9 range experts (top to bottom). Left: original label distribution colour-coded by range expert. Centre: original label vs. original model output. Right: original model output vs. output of the respective range experts. Note, that the results shown are from range experts retrained to obtain a more disentangled representation in connection with LRP.
Appendix D Augmenting wind turbine SCADA data with yaw-misalignment
We randomly add yaw misalignment of up to $15^{\circ}$ to our data SCADA set, and adjust the respective targets (turbine output) with a yaw misalignment factor $c_{ymis,i}=cos^{3}(\Delta_{yaw})$, if $v_{w,i}<v_{w,rated}$. This approximation can be easily derived from static flow equations and geometric considerations, for more details on how yaw misalignment affects turbine output see (Howland et al., 2020). After training and evaluation of the model on the augmented data, we can compare the magnitude of attributions to the ground truth.

### v3

@@ 0.C.2 Details quantitative evaluation
Figure 11: Overview of model performance on the rMNIST problem for 3, 6 and 9 range experts (top to bottom).
Figure 12: Overview model performance for the regression benchmarks (Sec. 5).
0.C.3 Augmenting wind turbine SCADA data with yaw-misalignment
We randomly add yaw misalignment of up to $15^{\circ}$ to our data SCADA set, and adjust the respective targets (turbine output) with a yaw misalignment factor $c_{ymis,i}=cos^{3}(\Delta_{yaw})$, if $v_{w,i}<v_{w,rated}$. This approximation can be easily derived from static flow equations and geometric considerations, for more details on how yaw misalignment affects turbine output see [17]. After training and evaluation of the model on the augmented data, we can compare the magnitude of attributions to the ground truth.
Figure 13: Overview model performance wind turbine example (Sec. 4.3)